# Uncertainty Propagation, and Parameter Calibration

### v1

@@ 3.1 Gaussian Processes (GPs)
Where $c_{k}$ is the output of the correlation function for the $k^{th}$ coordinate of the two input vectors. There are many options for correlation functions and here we use the rational quadratic kernel, which is created by adding squared exponential kernels across different length scales because it would capture some of the multiscale effects of the DNS. Other kernels failed, as discussed later. Overall, the unknown parameters in the GP formulation include the regression coefficients $\theta$, the variance $\sigma^{2}$, and the range parameters from the kernel as a vector $\bm{\gamma}$. In this formulation, the marginal likelihood function after integrating out ($\theta,\sigma^{2}$) becomes:
@@ 3.2 Hierarchical Multiscale Surrogate (HMS)
Figure 2: Behavior of the basis functions at increasing scales. Figure from [51].
Hierarchical and multiscale models typically have an identified hierarchy of approximations and make joint inference on the data by combining these model components in some rational fashion [52, 53]. Hierarchical models also have parallels to deep learning models which implement sequential function compositions to learn a data generation mechanism [54], multiresolution Analysis (MRA) implementing wavelet decompositions, ([55]), geometric analysis ([56, 57]) involving diffusion maps to analyze datasets at different scales and Hierarchical Radial Basis Functions (HRBF) ([58, 59, 60]). The combination of basis functions at different scales to produce a sparse representation is the core idea in most of these [52] though as we show in recent work [51], it is often efficient in the prediction phase to use only an optimal scale rather than a mix of scales. Hierarchical approximations also find applications in data visualization [61]. Recent work on explaining the success of deep learning has attributed it to the ability of these networks to accomplish hierarchical learning efficiently [62]. This approach for sparse representation, introduces a scale parameter $s$ and defines a mapping between $s$ and the corresponding approximation space $\mathcal{H}_{s}$ in a hierarchy of spaces. The idea of exploiting the inherent correlation structure in the data at multiple levels follows directly from [63, 64]. The hierarchical models are able to analyze irregularly structured data at different resolutions (scales), rapidly minimizing the error in fitting using scale and data dependent basis functions $B^{s}$, constructed by sampling suitable kernels at successively finer scales. Basis selection to approximate the observed data efficiently [40] uses a greedy scheme with forward selection and backward deletion phases that provides good performance avoiding the performance bottleneck of other strategies [51].
Figure 3: Hyperparameter tuning for the HMS surrogate. $\delta$ describes the upper bound tolerance for the current scale tolerance ($\epsilon_{s}<\delta$), $s$ is the number of scales, and $k=[1,5]$ refers to a training data fold. A vertical identifier is placed at the scale where the MSE is $<2e^{-05}$. Smaller compression ratios are better for fast prediction. Based on these results, we select $s=15,\delta=0.01$ for the two hyperparameters.
@@
Here $K_{s}=[b_{s}^{j}]$ is the full set of n functions at scale $s,p$ are some positive constant and $||\beta_{s}||_{0}$ counts the number of non-zero values in $\beta_{s}$. Since, this is a non-convex problem, an approximation of the problem corresponding to $||\cdot||_{1}$ norm is obtained. The greedy forward selection starts with an empty basis set $B_{s}$, and recursively adds functions from the set $K_{s}$ which provide the highest reduction in the current residual. This is continued until the tolerance $\epsilon_{s}$ is satisfied by the residual $r^{t}$ as $\frac{|r^{T}b^{j}_{s}|}{||b_{s}^{j}||_{2}^{2}}<\epsilon_{s}$. Before training the HMS surrogate, we performed hyperparameter tuning to select the number of scales and tolerance bounds. The tuning process is about the trade-off between approximation error and the compression ratio of the data $|C^{s}|/|X|$. As more scales $s$ are added, the training error reduces, while the need for more data in the approximation increases. The selected number of scales and tolerance are selected such that we achieve an acceptable order of error magnitude ($\sim e^{-05})$, while also maintaining the compression ratio as low as possible for fast prediction. Based on our test runs (see Fig. 3 for results from best runs), we selected $s=15$ scales and the tolerance bound as $\epsilon_{s}<\delta=0.01$.
@@ 3.3 Uncertain Inputs and Ensemble Strategy
The relevant input quantities to ABLATE are summarized in Table 1. To create accurate surrogates, the simulation ensemble needs to be selected such that it covers the domain and codomain of the simulation as best as possible. Some of the input quantities are considered uncertain and therefore have to be sampled for the ensemble. For the ensemble runs, the geometry remains fixed because the simulation runs for a relatively small real time (on average 25-75ms). We do not expect the geometry change to be significant during this short time given that the maximum regression rate of PMMA experimentally is measured up to 0.15mm/s [65], which justifies our assumption for the fixed geometry during a short simulation. We initially considered all possible reaction rate parameters of the Arrhenius expression (Eq. 11) and used local and global sensitivity analysis to select the reaction rate parameters to be included in the simulation ensembles, as described in Section 3.3.1. Only the selected reaction parameters are sampled for the ensembles, the remaining ones are fixed at expected values from literature. The inlet velocity factors are selected such that the ABLATE simulation corresponds to reasonable oxidizer flux $G$ of an equivalent slab burner experiment [5], so we selected velocities that correspond to $G{\in}[5,20]\frac{kg}{m^{2}s}$. Of the fuel parameters $\theta_{f}$, the latent heat of sublimation $l_{v}$ is considered uncertain and sampled given an estimated range from expert input, but is later calibrated with available experimental data. To generate the 64 simulation ensemble, the uncertain input quantities were selected with Latin Hypercube Sampling (LHS), where each sample is the only one in each axis-aligned hyperplane containing it. The values for the uncertain inputs of each simulation of the ensemble are shown in Table LABEL:tab:ensemble1 and LABEL:tab:ensemble2 in the Appendix.
@@
Velocity factor (m/s) $V_{in}$ Uncertain The inlet velocity profile is assumed parabolic
Activation Energies (cal/mol) $E_{*}$ Uncertain The energy required to form the transition state for a given reaction
@@
Pre-exponential factors (cm3 mol-1 cal -1) $A_{*}$ Uncertain Empirical constant between temperature and rate coefficient
@@ 3.3.1 Feature selection of Reaction Rate Parameters
    Feature selection between the reaction rate parameters with the objective being to minimize the testing MSE of the HMS surrogate. The other slab burner covariates ($x,y,l_{v},V_{in}$) all remain in the surrogate.
@@
where $T$ is the temperature, and $E_{i}$, $A_{i}$, and $\beta_{i}$ are the activation energy, pre-exponential factor, and temperature exponent coefficient of the $i$th reaction, respectively. By identifying the sensitive reactions via local sensitivity analysis, it becomes computationally feasible to perform VSA on their associated associated rate parameters ($A_{i},b_{i},E_{i}$). To conduct VGSA, we sample all parameters from a uniform probability distribution function (PDF) and compute the total Sobol sensitivity index ${S_{T}}_{i}=\frac{\mathbb{E}[Var(IDT_{i}|k_{\sim i})]}{Var(IDT)}$. However, in the combustion literature, it is known that there are multiple sources of correlation among these parameters [72] [73]. For our VSA, we consider a 20% uncertainty bound around the nominal values of the Arrhenius parameters associated with the most influential reactions and use TChem [74, 75, 76, 77, 78, 79] solvers for the sensitivity analyses. TChem computes the ignition delay time for a 0-dimensional constant volume homogeneous gas reactor. TChem defines IDT as the time when the build-up of $\mathrm{OH}$ radicals peaks. From the LSA, the most important reactions are the following:
@@
Figure 4 illustrates the results from the VGSA. The total sensitivity index is evaluated using 6000 samples drawn from the parameter distributions using LHS, using the Monte -Carlo (MC) estimator proposed by [80] and implemented in the parallel object-oriented library DAKOTA [81]. The figure shows that the activation energy in reaction 257 exhibits the most impact on ignition delay. This particular MMA reaction corresponds to chain branching highly reactive species, which play a crucial role in sustaining the combustion process by breaking down and producing radicals. Other important parameters include the activation energies of reaction 1 and 250 as well as the temperature exponent for reaction 249. Furthermore, we observed the rest of the parameters ($A250,b257,A1,E249,A249,A257$) to have a negligible effect on the IDT. Consequently, these parameters can be considered deterministic when producing simulation samples to train a surrogate model, as their uncertainties do not significantly impact the overall variability in the IDT.
Figure 4: Left: Effect of the most important Arrhenius parameters on ignition delay for our MMA mechanism. Right: Total sensitivity indices of the Arrhenius parameters of the four influential reactions identified by local sensitivity analysis. The 4 most important parameters ($E257,b249,E1,E250$) are considered most important and are therefore part of the uncertain simulation ensemble. The rest of the parameters are considered deterministic and are fixed, since they have minimal impact on ignition delay.
Figure 5: Surrogate feature selection between reaction rate parameters. The dashed red line indicates the minimum MSE. $E1$ and $b249$ achieve similar results. The test MSE is added over all testing folds. We note that adding all 4 of the reaction rate parameters results in an increase to testing MSE by an order of magnitude. Based on these results, we select $b249$ as the only reaction rate to be included in the surrogate.
To complete the feature selection, the testing MSE of the HMS surrogate is evaluated across different testing folds covering the entire dataset. Each testing fold is predicted with different models arising from the different combinations of reaction rate parameters and all the other slab burner predictor features ($x,y,l_{v},V_{in}$). The testing MSE evaluated here is an addition of the error at each fold. The results of the feature selection are shown in Fig. 5. Based on our analysis, we only include $b249$ as the only reaction rate feature in the surrogate models. While it may be unexpected that $b249$ alone produces the best surrogate model for generalization even if it is not the most sensitive parameter to ignition delay, we note that these parameters are not entirely independent of each other. Therefore, the sensitivity analysis (which is conducted with ignition delay as the dependent variable) is only a part of the selection process, whereas directly testing the surrogate model against the slab burner QoI (the regression rate) is ultimately the deciding factor, as regression rate is the quantity we want to predict.
@@ 3.4 Forward Uncertainty Propagation
After the development of the surrogates, they can be used to propagate forward the uncertainty in the inputs $\bm{x}=[x,y,b_{249},l_{v},V_{in}]$ to the QoI $\tilde{\dot{r}}$. The goal is to approximate the probability distribution of the regression rate and its statistics, which can be achieved via MC sampling methods [82] (Alg. 2). Forward UQ is possible given that the surrogates are fast to sample. To know when the MC sampling has converged, we monitor the mean and variance of the posterior distribution for the regression rate and stop when those its statistics difference are within a tolerance of $\epsilon_{*}=e^{-5}$. For the inputs, the velocity factor is assumed uniform throughout the parameter range. The fuel and chemistry parameters $l_{v},b_{249}$, are assumed normal with the mean being the nominal default DNS value and a variance that results in the distribution covering values ±$\pm 25\%$ from that nominal value. Given that the regression rate is highly dependent on the location of interest as known from our previous work with experiments, we consider the results in different regions of the slab burner. Experimentally, the front part ($x\leq 40$mm) is expected to achieve the highest regression due to the ramp effects and ignition location, the middle part ($40<x\leq 70$ mm) mostly achieves lower regression due to steady state combustion, and the rear part ($x>70$mm) also has higher regression due to re-circulation effects. In each case, we uniformly sample the possible boundary locations.
@@
$b_{249}\sim\mathcal{N}(2.54,0.029)(-)$
$l_{v}\sim\mathcal{N}(840890,4.91\times 10^{9})\frac{J}{kg}$
@@
Step 1. Draw a sample from each of the input quantities to create the new design vector $\bm{x^{*}}=[x,y,E^{*}_{1},l_{v}^{*},V_{in}^{*}]$
@@ 3.5 Model Calibration under Uncertainty
Finally, we demonstrate the Bayesian calibration of two key parameters: the latent heat of sublimation ($l_{v}$)–a fuel parameter in the DNS model—and the temperature exponent for reaction 249 ($b_{249}$)–a parameter in the chemical reaction mechanism. Both parameters are calibrated using regression rate measurements obtained from slab burner experiments. The calibration of physics-based models necessitates addressing uncertainties arising from incomplete and noisy experimental data, as well as the model’s limitations in capturing complex physical phenomena. Traditional deterministic inverse methods are insufficient for quantifying these uncertainties. Therefore, we adopt Bayesian approaches for statistical inverse analysis, as outlined in [83, 84, 85], which provide a probabilistic framework for UQ in model calibration. The solution to the Bayesian calibration problem is the probability distribution functions (PDFs) of the parameters ($\bm{x}_{calib}={l_{v},b_{249}}$) given the experimental data $\mathbf{D}$, such that
@@
Figure 6: Available regression rate measurements from an experiment with oxidizer flux $G=8.08\frac{kg}{m^{2}s}(\equiv V_{in}=6.258m/s)$.
@@ 4 Results and Discussion
This section provides the results after implementation of the methods introduced in Section 3. The section starts with a comparison via cross validation between the two surrogates GP and HMS. Then, we use HMS for forward UQ to estimate the probability distribution of the regression rate from the uncertain inputs in 3 regions of the slab burner: front, middle, and rear. Lastly, we demonstrate Bayesian inference of the fuel parameter latent heat of sublimation ($l_{v}$) and the temperature exponent of reaction 259 ($b_{249}$) with experimental slab burner regression rate data.
@@
To compare the two surrogate models, we first evaluate their training and testing accuracy on the available dataset via 5-fold cross validation [86]. We split the dataset of the 64 DNS simulations in 5 folds, 4 of which include the 13/64 simulations as part of the testing set and the remaining 51/64 simulations as training. The last fold includes 12/64 simulations as testing and the remaining 52/64 simulations as training. When a simulation is picked to be part of a training/testing set, all the boundary locations are included with it, we did not sample based on location. With the cross-validation approach, the goal is to evaluate the resulting surrogate models on the entire available dataset and particularly their ability to generalize well and predict correctly out-of-sample, i.e., from data they have not been trained on.
Fig. 7 shows some of the results from the first fold for both GP and HMS. Both surrogates are trained correctly and predict the training set perfectly. One obvious difference between the two models is that GPs inherently predict with a standard deviation which can be translated to a $95\%$ confidence interval, unlike HMS. Therefore, a GP may be more practical as evident in some of the results shown in Fig. 7 (e.g. see $IDs=8,12,13$), since it is able to capture all the simulation output within the confidence interval even if the mean is not precise. Comparing the GP mean with HMS, we conclude that both models appear to have similar performance out of sample, but the HMS follows the simulation data visibly closer when non-linear effects are present (e.g., see $IDs=2,11$). We also conclude that neither model is able to fully capture non-linearities as evident in e.g., $IDs=10,9,11$. Further investigation into this issue brought forward the following conclusions:
  1. 1.
    The GP performance is sensitive to the selection of kernel. Rational quadratic is the only kernel that produced a correct GP trained on the DNS data. The rational quadratic is a collection of exponential kernels, which captures some of the multiscale effects of the problem and justifies why the GP performs similarly to HMS. See Fig. 8(a) for an example of a poorly fit GP with Matérn kernel. The HMS on the other hand adaptively selects a set of basis functions that work for a large class of problems without heuristic choices of kernel.
  2. 2.
    Even though the DNS is solving a multiscale problem, both surrogate models are unable to capture all of the scales since the inputs are all far field inputs (far field velocity, parameters). For example, in Fig. 7, the non-linearities at the end of $ID=11$ are missed, and HMS is closer to the top of the regression rate spike than the GP. The models are attempting prediction of a boundary quantity that is dependent on the combustion process, but none of the far field inputs they are trained on fully carry this multiscale information. There is no inherent structure in the surrogates to be able to estimate the multiscale combustion process, unlike the DNS that fully solves the system.
  3. 3.
    As a demonstration, when a combustion-related quantity ($\dot{q}_{Net}$) is added as an input to the surrogate, the HMS fully captures the multiscale effects and is able to predict out of sample with precision as shown in Fig. 8(b). We note however that this would not be a useful surrogate in practice, as it is impossible to know the heat flux a priori without solving the combustion problem, which defeats the purpose of having a surrogate.
We show some more comparisons between the two models from the other cross-validation folds in the Appendix in Fig. LABEL:fig:HMSvGP_2.
@@
Fig. 9 shows the training and testing performance of the two models across all the data folds, measured as a normalized percentage error from the true value of the prediction. The GP, on average, is expected to achieve an error of $1.59\times 10^{-6}\%$ in the training set and $17.06\%$ when tested out of sample. The HMS, on average, is expected to achieve an error of $1.11\%$ in the training set and $16.62\%$ when tested out of sample. Overall, our conclusion is that both of these models are acceptable for usage in UQ or otherwise. The fact that both of these models are able to achieve an error $<20\%$ when predicting multiscale boundary quantities just from a few far field inputs is significant, particulary considering that they take a few $ms$ to run compared to 24 hours for the DNS.
@@
Figure 7: Comparison of GP surrogate and HMS for select simulation IDs belonging to the first fold of the cross validation. Figure titles mark whether the data were in the train or test set. As expected, both surrogates predict the training set perfectly. The GP inherently includes a standard deviation in the predictions which is translated to a $95\%$ confidence interval, adequately including the simulation data within these bounds. Both models seem to have similar mean predictions. When the regression rate includes nonlinear effects (e.g., spike at $70mm$ for $ID=2$ or spike at $80mm$ for $ID=11$ ), the HMS mean follows the trend visibly a bit closer than the GP.
(a) Fig. 8a     Example of a GP trained on the DNS data with Matérn kernel. The optimizer has not converged as evident by the very large variance bounds and the poor fit of the mean. Rational quadratic is the only kernel that produced a correct GP trained on the DNS data.
(b) Fig. 8b     Example of a HMS trained on the DNS data including the net radiation at the boundary ($\dot{q}_{Net}$). The HMS fully captures the multiscale effects and is able to predict out of sample with precision.
@@
Figure 9: Performance of GP and HMS across all data folds of the cross validation. The GP, on average, is expected to achieve an error of $1.59\times 10^{-6}\%$ in the training set and $17.06\%$ when tested out of sample. The HMS, on average, is expected to achieve an error of $1.11\%$ in the training set and $16.62\%$ when tested out of sample. Overall, our conclusion is that both of these models are acceptable for usage in UQ or otherwise.

### v2

@@ 2.1 Gaussian Processes (GPs)
Where $c_{k}$ is the output of the correlation function for the $k^{th}$ coordinate of the two input vectors. There are many options for correlation functions and here we use the rational quadratic kernel, which is created by adding squared exponential kernels across different length scales because it would capture some of the multiscale effects of the DNS. Overall, the unknown parameters in the GP formulation include the regression coefficients $\theta$, the variance $\sigma^{2}$, and the range parameters from the kernel as a vector $\bm{\gamma}$. In this formulation, the marginal likelihood function after integrating out ($\theta,\sigma^{2}$) becomes:
@@ 2.2 Hierarchical Multiscale Surrogate (HMS)
Figure 3: Behavior of the basis functions centered at the same point at increasing scales ($s$). Reprinted with permission from [52]. Copyright 2022 Elsevier under Creative Commons License.
Hierarchical and multiscale models typically have an identified hierarchy of approximations and make joint inference on the data by combining these model components in some rational fashion [53, 54]. HMS introduces a scale parameter $s$ and defines a mapping between $s$ and the corresponding approximation space $\mathcal{H}_{s}$ in a hierarchy of spaces. The idea of exploiting the inherent correlation structure in the data at multiple levels follows directly from [55, 56]. The hierarchical models are able to analyze irregularly structured data at different resolutions (scales), rapidly minimizing the error in fitting using scale and data dependent basis functions $B^{s}$, constructed by sampling suitable kernels at successively finer scales. Basis selection to approximate the observed data efficiently [38] uses a greedy scheme with forward selection and backward deletion phases that provides good performance avoiding the performance bottleneck of other strategies [52]. Fig. 3 shows the behavior of the basis functions centered at the same point at increasing scales $s$.
Figure 4: Hyperparameter tuning for the HMS surrogate. $\delta$ describes the upper bound tolerance for the current scale tolerance ($\epsilon_{s}<\delta$), $s$ is the number of scales, and $k$ refers to a training data fold. Based on these results, we select $s=21,\delta=0.005$ for the two hyperparameters.
@@
Here $K_{s}=[b_{s}^{j}]$ is the full set of n functions at scale $s,p$ are some positive constant and $||\beta_{s}||_{0}$ counts the number of non-zero values in $\beta_{s}$. Since, this is a non-convex problem, an approximation of the problem corresponding to $||\cdot||_{1}$ norm is obtained. The greedy forward selection starts with an empty basis set $B_{s}$, and recursively adds functions from the set $K_{s}$ which provide the highest reduction in the current residual. This is continued until the tolerance $\epsilon_{s}$ is satisfied by the residual $r^{t}$ as $\frac{|r^{T}b^{j}_{s}|}{||b_{s}^{j}||_{2}^{2}}<\epsilon_{s}$. Before training the HMS surrogate, we performed hyperparameter tuning to select the number of scales and tolerance bounds. The tuning process is about the trade-off between approximation error and the compression ratio of the data $|C^{s}|/|X|$. As more scales $s$ are added, the training error reduces, while the need for more data in the approximation increases. The selected number of scales and tolerance are selected such that we achieve an acceptable order of error magnitude (ł$\l e^{-05})$, while also maintaining the compression ratio as low as possible for fast prediction. Based on our test runs (see Fig. 4 for results from some runs), we selected $s=21$ scales and the tolerance bound as $\epsilon_{s}<\delta=0.005$ because this combination achieves the lowest MSE with a reasonable compression ratio. $\delta=0.001$ results in prohibitive compression ratios for a surrogate and $\delta=0.01$ would be acceptable as well, but it is preferred to achieve a lower MSE for only 0.1 increase in the compression ratio.
@@ 2.3 Uncertain Inputs and Ensemble Strategy
The relevant input quantities to ABLATE are summarized in Table 1. To create accurate surrogates, the simulation ensemble needs to be selected such that it covers the domain and codomain of the simulation as best as possible. Some of the input quantities are considered uncertain and therefore have to be sampled for the ensemble. For the ensemble runs, the geometry remains fixed because the simulation runs for a relatively small real time (on average 25-75ms). We do not expect the geometry change to be significant during this short time given that the maximum regression rate of PMMA experimentally is measured up to 0.15mm/s [57], which justifies our assumption for the fixed geometry during a short simulation. We initially considered all possible reaction rate parameters of the Arrhenius expression (Eq. 15) and used local and global sensitivity analysis to select the reaction rate parameters to be included in the simulation ensembles, as described in Section 3.3.1. Only the selected reaction parameters are sampled for the ensembles, the remaining ones are fixed at expected values from literature. The inlet velocity factors are selected such that the ABLATE simulation corresponds to reasonable oxidizer flux $G$ of an equivalent slab burner experiment [5], so we selected velocities that correspond to $G{\in}[5,20]\frac{kg}{m^{2}s}$. Of the fuel parameters $\theta_{f}$, the latent heat of sublimation $l_{v}$ is considered uncertain and sampled given an estimated range from expert input, but is later calibrated with available experimental data. To generate the 64 simulation ensemble, the uncertain input quantities were selected with Latin Hypercube Sampling (LHS), where each sample is the only one in each axis-aligned hyperplane containing it. The values for the uncertain inputs of each simulation of the ensemble are included in table format as supplemental material.
@@
Velocity factor (m/s) $V_{in}$ Uncertain Velocity profile is assumed parabolic
Activation Energies (cal/mol) $E_{*}$ Uncertain Energy required to form the transition state for a given reaction
@@
Pre-exponential factors (cm3 mol-1 cal -1) $A_{*}$ Uncertain Empirical constant associating temperature and rate coefficient
@@ 2.3.1 Feature selection of Reaction Rate Parameters
    Feature selection between the reaction rate parameters with the objective being to minimize the testing MSE of the HMS surrogate. The other slab burner covariates ($x,l_{v},V_{in}$) all remain in the surrogate.
@@
where $T$ is the temperature, and $E_{i}$, $A_{i}$, and $\beta_{i}$ are the activation energy, pre-exponential factor, and temperature exponent coefficient of the $i$th reaction, respectively. By identifying the sensitive reactions via local sensitivity analysis, it becomes computationally feasible to perform VSA on their associated associated rate parameters ($A_{i},b_{i},E_{i}$). To conduct VGSA, we sample all parameters from a uniform probability distribution function (PDF) and compute the total Sobol sensitivity index ${S_{T}}_{i}=\frac{\mathbb{E}[Var(IDT_{i}|k_{\sim i})]}{Var(IDT)}$. However, in the combustion literature, it is known that there are multiple sources of correlation among these parameters [64] [65]. For our VSA, we consider a 20% uncertainty bound around the nominal values of the Arrhenius parameters associated with the most influential reactions and use TChem [51] solvers for the sensitivity analyses. TChem computes the ignition delay time for a 0-dimensional constant volume homogeneous gas reactor. TChem defines IDT as the time when the build-up of $\mathrm{OH}$ radicals peaks. From the LSA, the most important reactions are the following:
@@
Figure 5 illustrates the results from the VGSA. The total sensitivity index is evaluated using 6000 samples drawn from the parameter distributions using LHS, using the Monte-Carlo (MC) estimator proposed by [66] and implemented in the parallel object-oriented library DAKOTA [67]. The figure shows that the activation energy in reaction 257 exhibits the most impact on ignition delay. This particular MMA reaction corresponds to chain branching highly reactive species, which play a crucial role in sustaining the combustion process by breaking down and producing radicals. Other important parameters include the activation energies of reaction 1 and 250 as well as the temperature exponent for reaction 249. Furthermore, we observed the rest of the parameters ($A_{250},b_{257},A_{1},E_{249},A_{249},A_{257}$) to have a negligible effect on the IDT. Consequently, these parameters can be considered deterministic when producing simulation samples to train a surrogate model, as their uncertainties do not significantly impact the overall variability in the IDT.
Figure 5: Left: Effect of the most important Arrhenius parameters on ignition delay for our MMA mechanism. Right: Total sensitivity indices of the Arrhenius parameters of the four influential reactions identified by local sensitivity analysis. The parameters ($E_{257},b_{249},E_{1},E_{250}$) are considered most important and are therefore part of the uncertain simulation ensemble.
Figure 6: Surrogate feature selection between reaction rate parameters. The dashed red line indicates the minimum MSE. $E_{1}$ and $(E_{1},b_{249})$ achieve similar results. The test MSE is added over all testing folds. Based on these results, we select $E1$ as the only reaction rate to be included in the surrogate.
To complete the feature selection, the testing MSE of the HMS surrogate is evaluated across different testing folds covering the entire dataset. Each testing fold is predicted with different models arising from the different combinations of reaction rate parameters and all the other slab burner predictor features ($x,l_{v},V_{in}$). The testing MSE evaluated here is an addition of the error at each fold. The results of the feature selection are shown in Fig. 6. Based on our analysis, we only include $E_{1}$ as the only reaction rate feature in the surrogate models. While it may be unexpected that $E_{1}$ alone produces the best surrogate model even if it is not the parameter with the highest sensitivity index, the sensitivity analysis (which is conducted with ignition delay as the dependent variable) is only a part of the selection process, whereas directly testing the surrogate model against the slab burner QoI (the regression rate) is ultimately the deciding factor, as regression rate is the quantity we want to predict.
@@ 2.4 Forward Uncertainty Propagation
After the development of the surrogates, they can be used to propagate forward the uncertainty in the inputs $\bm{x}=[x,E_{1},l_{v},V_{in}]$ to the QoI $\tilde{\dot{r}}$. The goal is to approximate the probability distribution of the regression rate and its statistics, which can be achieved via MC sampling methods [68] (Alg. 2). Forward UQ is possible given that the surrogates are fast to sample. To know when the MC sampling has converged, we monitor the mean and variance of the posterior distribution for the regression rate and stop when those its statistics difference are within a tolerance of $\epsilon_{*}=e^{-5}$. For the inputs, the velocity factor is assumed uniform throughout the parameter range. The fuel and chemistry parameters $l_{v},E_{1}$, are assumed normal with the mean being the nominal default DNS value and a variance that results in the distribution covering values ±$\pm 25\%$ from that nominal value. Given that the regression rate is highly dependent on the location of interest as known from our previous work with experiments, we consider the results in different regions of the slab burner. Experimentally, the front part ($x\leq 30$mm) is expected to achieve the highest regression due to the ramp effects and ignition location, the middle part ($30<x\leq 70$ mm) mostly achieves lower regression due to steady state combustion, and the rear part ($x>70$mm) also has higher regression due to re-circulation effects. In each case, we uniformly sample the possible boundary locations.
@@
$E_{1}\sim\mathcal{N}(15610,107.1)(cal/mol)$
$l_{v}\sim\mathcal{N}(840890,70074.16)\frac{J}{kg}$
@@
Step 1. Draw a sample from each of the input quantities to create the new design vector $\bm{x^{*}}=[x,E^{*}_{1},l_{v}^{*},V_{in}^{*}]$
@@ 2.5 Model Calibration under Uncertainty
Finally, we demonstrate the Bayesian calibration of two key parameters: the latent heat of sublimation ($l_{v}$)–a fuel parameter in the DNS model—and the activation energy for reaction 1 ($E_{1}$)–a parameter in the chemical reaction mechanism. Both parameters are calibrated using regression rate measurements obtained from slab burner experiments. The calibration of physics-based models necessitates addressing uncertainties arising from incomplete and noisy experimental data, as well as the model’s limitations in capturing complex physical phenomena. Traditional deterministic inverse methods are insufficient for quantifying these uncertainties. Therefore, we adopt Bayesian approaches for statistical inverse analysis, as outlined in [69, 70, 71], which provide a probabilistic framework for UQ in model calibration. The solution to the Bayesian calibration problem is the probability distribution functions (PDFs) of the parameters ($\bm{x}_{calib}={l_{v},E_{1}}$) given the experimental data $\mathbf{D}$, such that
@@
Figure 7: Available regression rate measurements from an experiment with oxidizer flux $G=8.08\frac{kg}{m^{2}s}(\equiv V_{in}=4.587m/s)$.
@@ 3 Results and Discussion
This section provides the results after implementation of the methods introduced in Section 3. The section starts with a comparison via cross validation between the two surrogates GP and HMS. Then, we use HMS for forward UQ to estimate the probability distribution of the regression rate from the uncertain inputs in 3 regions of the slab burner: front, middle, and rear. Lastly, we demonstrate Bayesian inference of the fuel parameter latent heat of sublimation ($l_{v}$) and the activation energy of reaction 1 ($E_{1}$) with experimental slab burner regression rate data.
@@
To compare the two surrogate models, we first evaluate their training and testing accuracy on the available dataset via 5-fold cross validation [72]. We split the dataset of the 64 DNS simulations in 5 folds, 4 of which include the 13/64 simulations as part of the testing set and the remaining 51/64 simulations as training. The last fold includes 12/64 simulations as testing and the remaining 52/64 simulations as training. When a simulation is picked to be part of a training/testing set, all the boundary locations are included with it, we did not sample based on location. With the cross-validation approach, the goal is to evaluate the resulting surrogate models on the entire available dataset and particularly their ability to generalize and predict correctly out-of-sample, i.e., from data they have not been trained on.
Figs. 8, 9 show some of the results from the first fold for both GP and HMS. Both surrogates predict the training set perfectly as expected. One obvious difference between the two models is that GPs inherently predict with a standard deviation which can be translated to a $95\%$ confidence interval, unlike HMS. For this particular problem, due to the multiscale nature of the data, the GP confidence intervals are extremely wide in all predictions. The wide bounds indicate low confidence in the model predictions, but can be useful in practice to identify an accurate bounded range for the regression rate since they cover most of the data. Comparing the GP mean with HMS, we conclude that both models appear to have similar performance out of sample for the portion of the slab after the initial non-linear peak of the regression rate. We also conclude that the GP is unable to predict the non-linear regression rate in the front of the slab for all IDs. The HMS predicts the non-linear part of the slab better as evident in IDs $=1,3,4,5$. In some cases (see IDs $=2,9,12$) the non-linear part at the front is not fully captured by the HMS, but the model form is able to follow the response, but miss the magnitude peak. We show some more comparisons between the two models from the other cross-validation folds in supplemental material. Overall, we conclude that the HMS due to its ability to capture multiscale effects is a better surrogate for accurate prediction as it is able to follow the regression rate response better out-of-sample. The information from the confidence interval of a GP is also useful in practice, as it provides a wide but valid range for the regression rate response.
@@
Fig. 10 shows the training and testing performance of the two models across all the data folds, measured as a normalized percentage error from the true value of the prediction. The GP, on average, is expected to achieve an error of $1.00\times 10^{-8}\%$ in the training set and $28.03\%$ when tested out of sample. The HMS, on average, is expected to achieve an error of $0.63\%$ in the training set and $13.37\%$ when tested out of sample. The HMS is robust in prediction throughout the entire dataset. Overall, our conclusion is that HMS is better than the GP for usage in UQ or otherwise for prediction in this multiscale problem. The fact that HMS is able to achieve an error $<15\%$ when predicting multiscale boundary quantities just from a few far field inputs is significant, particularly considering that it takes a few $ms$ to run compared to 24 hours for the DNS. Information from the GP is useful, if needing to bound the quantity within some range.
@@
Figure 8: Comparison of GP surrogate and HMS for select simulation IDs belonging to the first fold of the cross validation. Both surrogates predict the training set perfectly (shown in bounded box). When the regression rate includes nonlinear effects, the HMS mean follows the trend much closer than the GP and is overall more accurate.
Figure 9: Further comparison of GP surrogate and HMS for select simulation IDs belonging to the first fold of the cross validation. When the regression rate includes nonlinear effects, the HMS mean follows the trend much closer than the GP and is overall more accurate.
@@
Figure 10: Performance of GP and HMS across all data folds of the cross validation. The GP, on average, is expected to achieve an error of $1.00\times 10^{-8}\%$ in the training set and $28.03\%$ when tested out of sample. The HMS, on average, is expected to achieve an error of $0.63\%$ in the training set and $13.37\%$ when tested out of sample.